OpenStudy (anonymous):

Will give out medals!! Rational the denominator 9/sqrt x+3

OpenStudy (anonymous):

@rishavraj ?

OpenStudy (asad786):

rish are you muslim

rishavraj (rishavraj):

is it \[\frac{ 9 }{ \sqrt{x} + 3 }\]

OpenStudy (asad786):

just asking

OpenStudy (anonymous):

Yeah.

OpenStudy (asad786):

najia your muslim

OpenStudy (anonymous):

Yes, im muslim, does that matter though? :p I just need this problem solved.

OpenStudy (asad786):

@studygurl

OpenStudy (asad786):

srry

rishavraj (rishavraj):

@Asad786 no i m not...... @Najia2000 so it would be \[\frac{ 9 }{ \sqrt{x} + 3 } \times \frac{ \sqrt{x} - 3 }{ \sqrt{x} - 3 }\]

rishavraj (rishavraj):

SO IT WOULD BE \[\frac{ 9(\sqrt{x} - 3) }{ x - 9 }\]

OpenStudy (anonymous):

Thats the answer, but with plus sign. Thank you!

rishavraj (rishavraj):

-_- where plus sign???

rishavraj (rishavraj):

@Najia2000 u should hav told me before tht the question i wrote was wrong hav a look theres a difference ...in question its \[\sqrt{x + 3}\] bt i solved for \[\sqrt{x} + 3\]

OpenStudy (anonymous):

Oh, im reaaaalllly sorry. Sorry Sorry Sorry !! :/

rishavraj (rishavraj):

@Najia2000 u should be .... -_-

OpenStudy (anonymous):

I know, I am.

rishavraj (rishavraj):

heheheheh.......just kidding,......;P